13. Susie Sales signed a contract with a marketing firm for $12,500 per month for the next three years, payable at the end of the month. She also received a sign-on bonus of $25,000 that she received today. If she requires an annual return of 7.92%, what is the present value of this stream of cash flows?


14. Jeffrey wants to buy a $28,500 car. He is prepared to make a down payment of 10 percent of the purchase price and finance the remainder through his bank. The bank will loan him the money at an annual rate of 5.28 percent for a period of 60 months. What is the amount of the monthly payments?


15. Madeline would like to buy a delivery vehicle for her small business. She can afford to make a down payment of $7,500 and pay monthly (end-of-month) payments of $1,250. If the term is 48 months and the loan rate is 4.74 percent (APR), then what is the maximum that Madeline can pay for the delivery vehicle?
Answer #1

13.

Calculating Present Value,

PV = [FV = 0, T = 36, I = 0.0792/12, PMT = 12,500]

PV = 399,367.75

Present Value = 399,367.75 + 25,000

Present Value = $424,367.75

14.

Loan Amount = 0.90(28,500) = $25,560

Time Period = 60 months

Interest Rate = 5.28%

Calculating Monthly Payment,

PMT = [FV = 0, PV = 25,560, T = 60, I = 0.0528/12]

Monthly Payment = $485.63

15.

Calculating Present Value,

PV = [FV = 0, T = 48, PMT = 1,250, I = 0.0474]

PV = $54,557.27

Maximum Payment = 54,557.27 + 7,500

Maximum Payment = $62,057.27